OpenStudy (anonymous):

I will award you! :) Solve each system of equations by using elimination. 9p+8q=2 6p-5q=26 a. p = –0.67, q = 2.65 b. p = 2.34, q = –74 c. p = 2.34, q = –2.39 d. p = –1.9, q = –2.39

OpenStudy (anonymous):

multiply the first equation by 5 and the second equation by 8 5(9p+8q=2) 45p+40q=10 8(6p-5q=26) 48p-40q=208

OpenStudy (anonymous):

next eliminate the q's and combine like terms 45p+40q=10 48p-40=208 93p=218

OpenStudy (anonymous):

now divide 218 by 93 2.3

OpenStudy (anonymous):

2.34*

OpenStudy (anonymous):

that was p now lets plug p into the first equation

OpenStudy (anonymous):

9(2.34)+8q=2 21.06+8q=2

OpenStudy (anonymous):

subtract 21.06 to both sides 8q=-19.06 and divide by 8 q=-2.38

OpenStudy (anonymous):

understood? If not go here: http://openstudy.com/study#/updates/53cd6bade4b081ba9638ec9f

OpenStudy (anonymous):

@William.deems

OpenStudy (anonymous):

Thank you! :)